QUESTION 1 1 points Save Answer Which of the following about a wave pulse traveling down a long spring is true? i. The size of the wave pulse decreases significantly as the wave pulse travels a short distance i. The shape of the wave pulse gets spread out very quickly O Only i. O Only i. O Both i. and i. Neither i. nor ii. QUESTION 2 1 points Save Answer A wave pulse is described by a fuction y fx) att 0. If the wave pulse is traveling in the +x direction with a speed vO, which equation correctly represents the pulse at a later time t? QUESTION 3 1 points Save Answer How does increasing the tension in a string affect the speed of waves in the string? O The speed of the waves increases O The speed of the waves doesnt change The speed of the waves decreases
Answer #1

1). Correct option: neither i nor ii.

2). Correct option: y= f(x-vot)

3). Following the formula:

speed= \sqrt{\frac{Tension}{Mass. density}}

Correct option: The speed of the wave increases.